# Sergei Gukov, Ludmil Katzarkov, Josef Svoboda

### v2

@@ 4. Seifert manifolds
Let us fix some notation. [^9] We mostly follow [47] but use $b_{i}$ instead of $\omega_{i}$. In [30], the sign of $b$ and the role of $a_{i}$ and $b_{i}$ are flipped. Seifert (reduced) data consist of an integer $b$ and tuples $(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k})$ of integers such that $0<b_{k}<a_{k}$ and $\gcd(a_{k},b_{k})=1.$ Associated to it, there is a Seifert manifold $Y=M(b;(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k}))$ fibered over the sphere with $k$ singular fibers. We always assume that $k\geq 3$. Seifert manifolds with $k<3$ are lens spaces, and we can easily enlarge $k$ by adding tuples $(1,0)$. The manifold $Y$ can be described by a star-shaped plumbing graph with the central node having the framing $-b$ and with $k$ legs, see Fig. 6. The framing of vertices on $j$-th leg is given (starting at the node) by Hirzerbruch–Jung (HJ) continued fraction of $a_{j}/b_{j}$, namely the sequence $-c^{(j)}_{1},-c^{(j)}_{2},\dots,-c^{(j)}_{n_{j}}$ with $c^{(j)}_{1}\geq 1$ and $c^{(j)}_{i}\geq 2$ for $i=2,\dots,n_{j}$, such that
$\frac{a_{j}}{b_{j}}=[c^{(j)}_{1},c^{(j)}_{2},\dots,c^{(j)}_{n_{j}}]:=c^{(j)}_{1}-\cfrac{1}{c^{(j)}_{2}-\cfrac{1}{\ddots\,-\cfrac{1}{c^{(j)}_{n_{j}}}}}.$
Figure 6. Plumbing graph of a Seifert manifold.
We label the leaves of the plumbing graph as $v_{1},v_{2},\dots,v_{k}$ and the node as $v_{0}$. Denote $A=\prod_{i=1}^{k}a_{i}$, $A_{i}=A/a_{i}$ and $A_{ij}=A/{a_{i}a_{j}}$.
The fundamental group $\pi_{1}(Y)$ has a presentation given by elements $g_{1},g_{2},\dots,g_{k}$ and $g_{0}$ satisfying
$[g_{i},g_{0}]=1,g_{i}^{a_{i}}=g_{0},\text{ for }i=1,2,\dots,k,\quad\prod_{i=1}^{k}g_{i}^{b_{i}}=g_{0}^{b}$
The group $H=H_{1}(Y,\mathbb{Z})$ is its abelianization, and we denote the generators by the same letters. Note that $g_{i}$ can be identified with the classes $[e_{i}]\in H=\mathbb{Z}^{s}/M\mathbb{Z}^{s}$ of the canonical basis vectors for the corresponding leaf $v_{i}$ and $g_{0}=[e_{0}]$ is the class of the node.
Let
$e=-b+\sum_{i}\frac{b_{i}}{a_{i}}$
be the Euler number of the Seifert fibration. Seifert manifold $Y$ over $S^{2}$ can be represented by negative definite plumbing if and only if $e<0.$ Assuming this, $Y$ is a rational homology sphere and the order of the group $H=H_{1}(Y,\mathbb{Z})$ can be computed as:
$|H|=A|e|=|-Ab+\sum_{i}A_{i}b_{i}|.$
@@
As we saw in Section 3, some invariants of $Y$ can be read from the splice diagram of $Y$ while some need the full plumbing data. Our primary example are $Z_{0}(q)$ and $\widehat{Z}_{\sigma}(q)$, respectively. In the Seifert case, the splice diagram contains exactly the integers $a_{i}$; see Fig. 7, so this distinction is about (in)dependence on $b$ and $b_{i}$ [^10] And it is vacuous for integral homology spheres where $b$ and $b_{i}$ are uniquely determined by $a_{i}$..
One has different examples of this phenomenon: While for fixed $a_{1},a_{2},\dots a_{k}$, the order of $H$ can be arbitrarily large, the group $H/\langle g_{0}\rangle$ is independent of $b_{i}$, with order $A/\text{lcm}(a_{1},a_{2},\dots,a_{k}).$ Geometrically, $g_{0}$ is the generator corresponding to the fiber of the Seifert fibration, and $H/\langle g_{0}\rangle$ is the fundamental group of the base orbifold $S^{2}$. Another interesting example is provided by the counts of $\mathop{SL}(2,\mathbb{C})$ vs $\mathop{SU}(2)$ connections, see Section 5.
Figure 7. Splice diagram of Seifert manifold $M(b;(a_{1},b_{1}),\dots,(a_{k},b_{k}))$.
4.1. Reduction theorem for Seifert manifolds
@@
The main theorem of this section, Theorem 4.2, generalizes these results to any number of fibers, removes the conditions on $H$ and, perhaps more importantly, makes the role of Seifert data more transparent. It also emphasizes the role of the series $Z_{0}(q)$, treating $\widehat{Z}_{\sigma}(q)$ as its $H$-equivariant refinement, using the anticanonical $spin^{c}$ structure $\sigma^{\text{ac}}$ from Section 2.1. This “reduction theorem” gives us a formula for $Z_{0}(q)$ and $\widehat{Z}_{\sigma}(q)$ using single variable rational functions. Recall the “Laplace transform” $\mathcal{L}_{A}$ defined as
(4.1) $\mathcal{L}_{A}(t^{n})=q^{n^{2}/4A}.$
@@
Let $Y=M(b;(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k}))$ be a Seifert manifold over $S^{2}$ ($k\geq 3$) with $e<0$. Then
(4.2) $Z_{0}(q^{|H|})=q^{\Lambda}\mathcal{L}_{A}(\operatorname{s.e.}f_{0}(t)),$

### v3

@@ 4 Seifert manifolds
Let us fix some notation. We mostly follow [47] but use $b_{i}$ instead of $\omega_{i}$. In [29], the sign of $b$ and the roles of $a_{i}$ and $b_{i}$ are flipped. Seifert (reduced) data consist of an integer $b$ and tuples $(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k})$ of integers such that $0<b_{i}<a_{i}$ and $\gcd(a_{i},b_{i})=1$ for $i=1,2,\dots,k$. Associated with it is a Seifert manifold $Y=M(b;(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k}))$ fibered over the sphere with $k$ singular fibers. We always assume that $k\geq 3$. Seifert manifolds with $k<3$ are lens spaces, and we can easily enlarge $k$ by adding tuples $(1,0)$. The manifold $Y$ can be described by a star-shaped plumbing graph with the central node having the framing b-b and with $k$ legs, see Figure 6. The framing of vertices on $j$-th leg is given (starting at the node) by Hirzerbruch–Jung (HJ) continued fraction of $a_{j}/b_{j}$, namely the sequence $c^{(j)}_{1},-c^{(j)}_{2},\dots,-c^{(j)}_{n_{j}}$ with $c^{(j)}_{1}\geq 1$ and $c^{(j)}_{i}\geq 2$ for $i=2,\dots,n_{j}$, such that
$\frac{a_{j}}{b_{j}}=\big{[}c^{(j)}_{1},c^{(j)}_{2},\dots,c^{(j)}_{n_{j}}\big{]}:=c^{(j)}_{1}-\cfrac{1}{c^{(j)}_{2}-\cfrac{1}{\ddots-\cfrac{1}{c^{(j)}_{n_{j}}}}}$.
Figure 6: Plumbing graph of a Seifert manifold.
We label the leaves of the plumbing graph as $v_{1},v_{2},\dots,v_{k}$ and the node as $v_{0}$. We denote
$A=\prod_{i=1}^{k}a_{i},\qquad A_{i}=\frac{A}{a_{i}},\qquad A_{ij}=\frac{A}{a_{i}a_{j}}.$
The fundamental group $\pi_{1}(Y)$ has a presentation given by elements $g_{1},g_{2},\dots,g_{k}$ and $g_{0}$ satisfying $g_{i},g_{0}]=1$, $g_{i}^{a_{i}}=g_{0}$, for $i=1,2,\dots,k$, $\prod_{i=1}^{k}g_{i}^{b_{i}}=g_{0}^{b}$. The group $H=H_{1}(Y,\mathbb{Z})$ is its abelianization, and we denote the generators by the same letters. Note that $g_{i}$ can be identified with the classes $e_{i}]\in H=\mathbb{Z}^{s}/M\mathbb{Z}^{s}$ of the canonical basis vectors for the corresponding leaf $v_{i}$ and $g_{0}=[e_{0}]$ is the class of the node.
Let $e=-b+\sum_{i}\frac{b_{i}}{a_{i}}$ be the Euler number of the Seifert fibration. Seifert manifold $Y$ over $S^{2}$ can be represented by a negative definite plumbing if and only if $e<0$. Assuming this, $Y$ is a rational homology sphere and the order of the group $H=H_{1}(Y,\mathbb{Z})$ can be computed as
$|H|=A|e|=\biggl{|}-Ab+\sum_{i}A_{i}b_{i}\biggr{|}.$
@@
As we saw in Section 3, some invariants of $Y$ can be read from the splice diagram of $Y$ while others need the full plumbing data. Our primary example are $Z_{0}(q)$ and $\widehat{Z}_{\sigma}(q)$, respectively. In the Seifert case, the splice diagram contains exactly the integers $a_{i}$; see Figure 7, so this distinction is about (in)dependence on $b$ and $b_{i}$. [^11] And it is vacuous for integral homology spheres where $b$ and $b_{i}$ are uniquely determined by $a_{i}$.
One has different examples of this phenomenon: While for fixed $a_{1},a_{2},\dots,a_{k}$, the order of $H$ can be arbitrarily large, the group $H/\langle g_{0}\rangle$ is independent of $b_{i}$, with order $A/{\rm lcm}(a_{1},a_{2},\dots,a_{k})$. Geometrically, $g_{0}$ is the generator corresponding to the fiber of the Seifert fibration, and $H/\langle g_{0}\rangle$ is the fundamental group of the base orbifold $S^{2}$. Another interesting example is provided by the counts of $\operatorname{SL}_{2}(\mathbb{C})$ vs $\operatorname{SU}(2)$ connections, see Section 5.
Figure 7: Splice diagram of Seifert manifold $M(b;(a_{1},b_{1}),\dots,(a_{k},b_{k}))$.
4.1 Reduction theorem for Seifert manifolds
@@
The main theorem of this section, Theorem 4.2, generalizes these results to any number of fibers, removes the conditions on $H$ and, perhaps more importantly, makes the role of Seifert data more transparent. It also emphasizes the role of the series $Z_{0}(q)$, treating $\widehat{Z}_{\sigma}(q)$ as its $H$-equivariant refinement, using the anticanonical ${\rm spin}^{c}$ structure $\sigma^{{\rm ac}}$ from Section 2.1. This “reduction theorem” gives us a formula for $Z_{0}(q)$ and $\widehat{Z}_{\sigma}(q)$ using single variable rational functions. Recall the “Laplace transform” $\mathcal{L}_{A}$ defined as
$\mathcal{L}_{A}(t^{n})=q^{n^{2}/4A}.$ (4.1)
@@
Let $Y=M(b;(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k}))$ be a Seifert manifold over $S^{2}$ $(k\geq 3)$ with $e<0$. Then $Z_{0}\bigl{(}q^{|H|}\bigr{)}=q^{\Lambda}\mathcal{L}_{A}(\mathrm{s.e.}\,f_{0}(t))$, where
$f_{0}(t)=\frac{\bigl{(}t^{A_{1}}-t^{-A_{1}}\bigr{)}\bigl{(}t^{A_{2}}-t^{-A_{2}}\bigr{)}\cdots\bigl{(}t^{A_{k}}-t^{-A_{k}}\bigr{)}}{\bigl{(}t^{A}-t^{-A}\bigr{)}^{k-2}},$ (4.2)
$\widehat{Z}_{h\sigma^{{\rm ac}}}(q)=q^{\Lambda/|H|}\mathcal{L}_{A|H|}(\mathrm{s.e.}\,f_{h\sigma^{{\rm ac}}}(t)),$